OpenStudy (anonymous):

the volume of a rectangular chinese box is given by the formula v=lwh. the volume of the box shown on the right is 480 cu units. find the width and length. (height is 6) length is x+2 (width is x) the width of the box is () units the length of the box is () units

OpenStudy (anonymous):

(x+2)(x)(6)=480

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

\[x^2+2x = 80\]\[x^2+2x-80=0\]\[(x+10)(x-8)=0\] so, x = 8 Length: 10 Width: 8 Height: 6

OpenStudy (anonymous):

it can't be -10 cuz thats stupid